OpenStudy (anonymous):

to get rid of the radical in the problem, the square root of 2 times q equals p, i understand that you square both sides but do you square both sides bc its 2? or do you square both sides bc its the square root?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

because its square root

OpenStudy (anonymous):

\[\sqrt{2}\sqrt{2}=2\]

OpenStudy (anonymous):

thank you! :) ok so in the problem 2q^2 =p^2 in what way does 2 divide the left hand side?

OpenStudy (anonymous):

Not sure I understand your question. However, if you wanted to isolate and solve for q. You would begin by dividing both sides by 2 (what you do to one side of an equation, you must also do to the other.)\[(2/2)q ^{2}=p ^{2}/2\]\[q ^{2}=p ^{2}/2\]